OpenStudy (anonymous):

Six pears and three apples cost $3.90. Two pears and five apples cost $3.30. How much does one pear cost?

OpenStudy (texaschic101):

6p + 3a = 3.90 2p + 5a = 3.30 Those are you 2 equations. Are you familiar with the elimination method ?

OpenStudy (anonymous):

not really can u help me

OpenStudy (anonymous):

wouldnt it be .30 cents

OpenStudy (texaschic101):

6p + 3a = 3.90 -->(5) 2p + 5a = 3.30 -->(-3) ------------------ 30p + 15a = 19.50 (result of multiplying by 5) -6p - 15a = -9.9 (result of multiplying by -3) ------------------add 24p = 9.6 p = 9.6/24 p = .40 cents <-- your answer Do you have any questions ?

OpenStudy (anonymous):

lol i was sooo close thank u :)

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

im sorry wouldnt it be 12 for 6p*2p i was confused u got 30p

OpenStudy (texaschic101):

no, because I multiplied the first equation by 5 and the second equation by -3. I did this so that the a's would cancel out...be eliminated, hence the elimination method.

OpenStudy (anonymous):

ooo ok i got confused lol sorry

OpenStudy (anonymous):

i got more

OpenStudy (texaschic101):

lf you are so unsure, I can finish the problem and solve for a, then sub our known values in and check the problem...would you feel better if I did that ?

OpenStudy (anonymous):

yes plzzz

OpenStudy (texaschic101):

ok...we know that p = .40 (40 cents)...lets sub .40 in for p in either of the original equations. 2p + 5a = 3.30 2(.40) + 5a = 3.30 .80 + 5a = 3.30 5a = 3.30 - .80 5a = 2.50 a = 2.50/5 a = .50 (50 cents) Now we will check our answers... 6p + 3a = 3.90 6(.40) + 3(.50) = 3.90 2.40 + 1.50 = 3.90 3.90 = 3.90 (correct) a = .50 (50 cents) and p = .40(40 cents)

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (texaschic101):

I understand that you have more problems to do ? Do you need help ?

OpenStudy (anonymous):

yes i do thank you soo much :)

OpenStudy (texaschic101):

are your questions posted or do you want to ask them here ?